Simplify the expression. Assume all variables are positive. Leave answer in radical form, no decimals. Let f(x)=16-x^2 and g(x)=4-x. Find f(x)-g(x)
Answer by edjones(8007)

Your question isn't clear. There are no radicals.
f(x)=16-x^2
g(x)=4-x
f(x)-g(x)
=16-x^2-(4-x)
=16-x^2-4+x
=-x^2+x+12
=-(x^2-x-12)
=-(x-4)(x+3)
.
Ed